# Constraint-Driven Multi-USV Coverage Path Generation for Aquatic Environmental Monitoring

### v2

@@ Abstract
In this article, we address aquatic environmental monitoring using a fleet of unmanned surface vehicles (USVs). Specifically, we develop an online path generator that provides either of circular or elliptic paths based on the real-time feedback so that the USVs efficiently sample the sensor data over given aquatic environment. To this end, we begin by formulating a novel online path generation problem for a group of Dubins vehicles in the form of cost minimization based on the formulation of persistent coverage control. We then transform the cost minimization into a constraint-based specification so that a prescribed performance level is certified. An online coverage path generator is then designed based on the so-called constraint-based control in order to meet the performance certificate together with additional constraints inherent in the parameters that specify the paths. It is also shown there that the present constraint-based approach allows one to drastically reduce the computational complexity stemming from combinations of binary variables corresponding to the turning directions of the USVs. The present coverage path generator is finally demonstrated through simulations and experiments on an original testbed of multiple USVs.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
From methods available in the literature, we are especially interested in coverage control, which distributes a group of mobile sensors over the environment for efficient data sampling by following the gradient over an objective function [11]. From this basic setup, some variations have been proposed. For example, persistent coverage control stimulates the mobile sensors to continuously patrol the environment rather than forming a stationary configuration [14, 15, 16]. Other recent advances in coverage control include data-driven policies and dealing with unknown event density functions [17, 18, 19]. Finally, a significant topic in this context is that of constraints, which can also be accounted for by means of constraint-based control [20]. In this way, the coverage control policy can be aware of issues such as the battery levels —essential the for long-duration autonomy of the robots [21]—, obstacles, and performance guarantees [15, 22, 23].
@@
Additionally, the kinematic vehicle models assumed in the above papers allow instantaneous velocity changes, but USVs may find problems to follow the velocities generated by coverage controllers due to their greater inertia. Another issue is that water vehicles may suffer significant disturbances due to currents, waves, and wind. For these reasons, in marine craft control it is common to employ a hierarchical architecture composed of control layers for guidance, navigation, and control [30]. Indeed, multiple works follow a backstepping approach [31, 32], where the overall control problem is decoupled into a high-level problem that generates ideal paths to follow and a low-level problem that makes USVs follow the path while rejecting disturbances. That is, the fleet trajectory design is decoupled from the individual trajectory tracking problem, as happens, for example, in [33]. With this in mind, it is not surprising that most results on coverage control with USVs focus on coverage path planning [34, 35, 36, 37], leaving path following to low-level controllers. Nevertheless, even in these simpler setups, there are challenges to solve. For example, it is difficult to update the paths flexibly under real situations because path planning is executed in a different layer. In addition, the path-planning approach generates only a finite-length path, which may not be suitable for persistent monitoring of the aquatic environment. Note that the concept of the constraint-based control has been recently applied to path planning or trajectory generation e.g. in [38, 39, 40, 41, 42]. However, the above shortcomings of [34, 35, 36, 37] apply to the papers [38, 39, 40, 41], and all of [38, 39, 41, 42] focus on safety certificates and cannot be a solution to the area coverage problem.
@@
    We apply the constraint-based control to the high level path generation, merging path generation and path following layers by exploiting a suitable trajectory parametrization for the nature of the robots considered. This can provide advantages because it combines the best of the two layers: coverage-performance informed decision making and real time feedback.
@@ II Problem Formulation
The position coordinates of the origin and rotation angle of $\Sigma_{i}$ relative to the world frame $\Sigma_{\mathrm{w}}$ are denoted by $p_{i}\in\mathbb{R}^{2}$ and $\theta_{i}\in\mathbb{S}^{1}$, respectively. We suppose that the motion of USV $i$ obeys so-called Dubins vehicle model
$\begin{bmatrix}\dot{p}_{i}\\ \dot{\theta}_{i}\end{bmatrix}=\begin{bmatrix}\cos\theta_{i}&0\\ \sin\theta_{i}&0\\ 0&1\end{bmatrix}\nu_{i},\ \nu_{i}=\begin{bmatrix}\bar{v}\\ \omega_{i}\end{bmatrix},$ (1)
@@
where $\overline{\delta}$, $\underline{\delta}$ are positive scalars and the operator
@@
In this article, we design an online path generator to efficiently sample data on all points $q_{j}\ (j=1,2,\dots,m)$. In principle, our solution can be applied to any path that meet the following conditions:
@@
Note that the present solution updates the path and hence the actual trajectory that each USV follows can be more complex than the circular or elliptic path.
@@ II-A Circular Path Generation Problem
respectively. Solving the quadratic equation (5a) together with (5b) for $c_{i}$ yields two solutions, which correspond to the right-hand turning path, denoted by $\mathrm{r}$, and left-hand one, denoted by $\mathrm{l}$, as illustrated in Fig. 2. Thus, the turning direction becomes another decision variable for each USV $i$, which is denoted by ${\rm X}_{i}\in\mathcal{X}=\{\mathrm{r},\mathrm{l}\}$. Notice that once the turning direction ${\rm X}_{i}\in\mathcal{X}$ is fixed, we can eliminate the variable $c_{i}$ by solving (5) in advance. Specifically, given $z_{i}(t)=[p_{i}^{\top}(t)\ \theta_{i}(t)]^{\top}$, the center $c_{i}$ meeting (5) is given by
@@
following the manner of coverage control [11, 14, 15, 16], where $c=[c^{\top}_{1}\ c^{\top}_{2}\ \cdots\ c^{\top}_{n}]^{\top}$, $r=[r_{1}\ r_{2}\ \cdots\ r_{n}]^{\top}$, and $\mathrm{X}=(\mathrm{X}_{i})_{i=1}^{n}$. Now, since both of the function $J_{\rm c}^{\mathrm{X}}$ and the constraints (5) are changing in time due to the variations of $\phi_{j}(t)$ and $z_{i}(t)$ in real time, it is difficult to maximize $J_{\rm c}^{\mathrm{X}}$ while meeting (5) instantaneously. Also, the gradient-based methodology in [14] does not provide any performance guarantee. We thus relax the objective and consider the constraint-based specification
@@ III-B Elliptic Coverage Path Generator
In order to bound the size of the ellipse, we constrain the eigenvalues of the matrix $S_{i}$ as
@@
Especially for large USV fleets, no observation point $q_{j}$ may be assigned to a USV as a result of partitioning in (28) or (41). When this happens, all the coefficients for the variable $\rho_{i}$ in the inequality constraint on the performance guarantee in (33) or (46) get equal to zero, and hence the constraint gets independent of the selection of $\rho_{i}$. Noticing that the cost in the QP is given by $\|\rho_{i}\|^{2}$ plus the cost for the slack variable, the optimal solution must be $\rho_{i}=0$ and hence the previous path parameters must be maintained. Thus, even when no observation point is assigned to a USV, no serious problem happens except for the violation of the local performance constraint.
@@
The Voronoi-like partition may have instantaneous big changes and accordingly the paths may significantly change, especially for large USV fleets. Even in the presence of such changes, the paths must continue to meet (5) or (15) as well as the size constraints (27) or (40). Consequently, such changes do not impose extreme actions on the USVs. In other words, the effect of possible instability of the Voronoi-based allocation is inherently mitigated by these constraints.
@@ IV Simulation with Ideal Mathematical Model
Figure 7: Time evolution of the functions $b^{2}_{\mathrm{e},1}$ (cyan), $b^{3}_{\mathrm{e},1}$ (orange), $b^{4}_{\mathrm{e},1}$ (yellow), $b^{5}_{\mathrm{e},1}$ (purple) associated with the elliptic shape and size constraints for USV 1.
The movie of the simulation is uploaded to https://youtu.be/DfWsMAWdaZc?si=3H2nfHVnl7wsEAKP, whose snapshots are shown in Fig. 5. The white curves show the trajectories of the USVs. Since the paths are updated online, the trajectories are no longer the ellipses. The color map of the environment shows the value of the importance index $\phi_{j}$, where red region indicates high values and blue low values. We observe from the movie and the snapshots that both USVs update the elliptic paths so that they pass through the red area.
The time evolution of the functions $b^{1}_{\mathrm{e},1}$ (red) and $b^{1}_{\mathrm{e},2}$ (blue) are shown in Fig. 7. We observe from the figure that $b^{1}_{\mathrm{e},i}\geq 0$ is satisfied at almost all times. This implies that the present path generator almost certifies the performance guarantee. However, the constraint is violated only at around $160$\mathrm{s}$$. In order to further investigate the violation, we show the time evolution of the functions $b^{k}_{\mathrm{e},1}\ (k=2,3,4,5)$ for USV 1 in Fig. 7. We see from this figure that $b^{5}_{\mathrm{e},1}$ is close to the lower limit and the constraint associated with $b^{5}_{\mathrm{e},1}$ gets active during the period. We conclude from these figures that the constraint violations in Fig. 7 are caused by the relaxation of the performance constraints to soft constraints in (46). Note that the constraint on $b^{5}_{\mathrm{e},1}$ is not only violated by the initial state but also slightly violated at around 160s. The latter violation was caused by the unavoidable discretization in the path computation. It is however well-known that the constraint-based control inherently drives the states to satisfy the constraints even if the state gets out of the safe set for some reason. This is why these temporal constraint violations do not cause any serious problem.
From these results, we conclude that the present path generator works as expected, at least for an ideal mathematical model (1). Finally, in order to show that the proposed path generator works even for a larger fleet of USVs, a simulation video with 6 USVs is uploaded to https://youtu.be/oS3x3O_IQB8 for reference.
@@ V-B Modeling and Local Controller Design
Figure 12: Time series data of the model outputs (blue), and the actual output data (gray) and the input data (black).
@@
where the gains were tuned so that the gain crossover frequency was equal to 4.0 rad/s. A bigger crossover frequency was prohibited due to the actuator saturation.  [^4] The actuators suffer from saturation in order to prevent the thrusters from rotating backwards. Specifically, both of $u_{i}^{\rm right}$ and $u_{i}^{\rm left}$ in (50) should remain greater than 1500. The time response for the step reference $\omega_{i}^{\rm ref}=0.5$\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$$ is shown in Fig. 13. We see that $\omega_{i}$ successfully tracks to the reference $\omega_{i}^{\rm ref}$ despite the disturbances from the water.
@@ V-C Collision Avoidance against Walls
(a) $t=5$\mathrm{s}$$
(b) $t=30$\mathrm{s}$$
(c) $t=60$\mathrm{s}$$
(d) $t=120$\mathrm{s}$$
(e) $t=180$\mathrm{s}$$
(f) $t=240$\mathrm{s}$$
Figure 15: Snapshots of the experiment at each time. The left side shows moving two real USVs on the experiment pool. The red and green circular paths on the right side, indicate paths corresponding to USVs 1 and 2. White curves show the trajectories of the two USVs after $t=0$\mathrm{s}$$.
@@ V-D Experimental Verification
We implemented the proposed control architecture for two USVs under the settings above. The movie of the experiment is uploaded to https://youtu.be/YuZbbcMRrLU?si=PAqEAuRi-jN2nSwq, whose snapshots are shown in Fig. 15. We see from the movie and figures that the USVs switch the turning directions $\mathrm{X}_{i}$ and radius $r_{i}$ so that they are persistently driven to the red regions with high importance indices, while avoiding overlaps of the monitoring areas between the two USVs.
@@
Let us finally show the actual monitoring performance achieved by the USVs. To this end, we show in Fig. 18 the time evolution of the sum of the overall importance indices, $\sum_{j=1}^{m}\phi_{j}$, which is regarded as a reasonable metric for the monitoring performance. It is seen from this figure that it initially decays sharply and then is maintained below $2.1\times 10^{3}$. The fact that the performance metric stays at a certain value means that each observation point has been monitored by the USVs persistently. It is thus concluded that the present coverage path generator successfully achieves persistent environmental monitoring even in the presence of uncertainties in the physical world. Note however that the relation between the performance level $\gamma$ for the path generation and the actual environmental monitoring performance remains implicit. A direct evaluation of the latter is left as future work.
@@ V-E Comparative Evaluation with Baseline Algorithm
These paths are discretized by $0.2$\mathrm{m}$$ and the waypoints $\varpi_{i}^{1},\varpi_{i}^{2},\cdots,\varpi_{i}^{\ell}$ are assigned at the endpoints of the path segments, with $\varpi_{i}^{\ell+1}=\varpi_{i}^{1}$. The closest segment is then selected as the initial target path segment for a path following controller based on the Line-of-Sight guidance algorithm [30]. Given the target path segment $\varpi_{i}^{k}\varpi_{i}^{k+1}$, the reference angle for USV $i$ is determined as follows:
@@
The video of the baseline algorithm experiment is available at https://youtu.be/ACtlfoO1MlU?si=uVJx-2Ql_PtJFcsM. The path is set up so that USVs do not collide with the pool wall. This is especially the case when the USV follows a path segment connecting the end and beginning points of the path generated by the DCAC algorithm. It takes about $55$\mathrm{s}$$ for a USV to complete one lap. When the USV follows the straight line part, the observation area tends to cover more points with higher importance indices. When it follows the curves, it tends to cover points that have been recently observed. The comparison of the sum of importance indices between the baseline algorithm and the proposed method is shown in Fig. 20. In both cases, the USVs pass through unvisited points at the beginning, so that there is no significant difference in the first $10$\mathrm{s}$$. The steep declines continue in the first $50$\mathrm{s}$$ in both methods but the coverage path generation achieves lower values than the lawnmower pattern. After the rapid decline, the values are stabilized, albeit with some oscillations. The results demonstrate that our proposed method achieves more efficient coverage by reducing the sum of importance indices more effectively than the baseline algorithm.
In addition, the present method is advantageous over the conventional path planning methods in the following aspects. First, the proposed path generator explicitly reflects the decay/increase rate of the importance indices which correspond to the required amount of the sensing data and the required timing of resampling. Second, the generator can compensate for the coverage holes caused e.g. by disturbances due to the real-time feedback of the current importance indices. Third, the generator has flexibility in adding or deleting USVs in the middle of the operation, since the area partitions are updated in real-time.

### v3

@@ Abstract
In this article, we address aquatic environmental monitoring using a fleet of unmanned surface vehicles (USVs). Specifically, we develop an online path generator that provides either circular or elliptic paths based on the real-time feedback so that the USVs efficiently sample the sensor data over given aquatic environment. To this end, we begin by formulating a novel online path generation problem for a group of Dubins vehicles in the form of cost minimization based on the formulation of persistent coverage control. We then transform the cost minimization into a constraint-based specification so that a prescribed performance level is certified. An online coverage path generator is then designed based on the so-called constraint-based control in order to meet the performance certificate together with additional constraints inherent in the parameters that specify the paths. It is also shown there that the present constraint-based approach allows one to drastically reduce the computational complexity stemming from combinations of binary variables corresponding to the turning directions of the USVs. The present coverage path generator is finally demonstrated through simulations and experiments on an original testbed of multiple USVs.
@@
This work has been accepted for publication in IEEE Transactions on Control Systems Technology. © 2025 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission. This is the author’s version which has not been fully edited and content may change prior to final publication. Citation information: DOI 10.1109/TCST.2025.3572776
@@ I Introduction
From methods available in the literature, we are especially interested in coverage control, which distributes a group of mobile sensors over the environment for efficient data sampling by following the gradient over an objective function [11]. From this basic setup, some variations have been proposed. For example, persistent coverage control stimulates the mobile sensors to continuously patrol the environment rather than forming a stationary configuration [14, 15, 16]. Other recent advances in coverage control include data-driven policies and dealing with unknown event density functions [17, 18, 19]. Finally, a significant topic in this context is that of constraints, which can also be accounted for by means of constraint-based control [20]. In this way, the coverage control policy can be aware of issues such as the battery levels —essential for the long-duration autonomy of the robots [21]—, obstacles, and performance guarantees [15, 22, 23].
@@
Additionally, the kinematic vehicle models assumed in the above papers allow instantaneous velocity changes, but USVs may find problems to follow the velocities generated by coverage controllers due to their greater inertia. Another issue is that water vehicles may suffer significant disturbances due to currents, waves, and wind. For these reasons, in marine craft control, it is common to employ a hierarchical architecture composed of control layers for guidance, navigation, and control [30]. Indeed, multiple works follow a backstepping approach [31, 32], where the overall control problem is decoupled into a high-level problem that generates ideal paths to follow and a low-level problem that makes USVs follow the path while rejecting disturbances. That is, the fleet trajectory design is decoupled from the individual trajectory tracking problem, as happens, for example, in [33]. With this in mind, it is not surprising that most results on coverage control with USVs focus on coverage path planning [34, 35, 36, 37], leaving path following to low-level controllers. Nevertheless, even in these simpler setups, there are challenges to solve. For example, it is difficult to update the paths flexibly under real situations because path planning is executed in a different layer. In addition, the path-planning approach generates only a finite-length path, which may not be suitable for persistent monitoring of the aquatic environment. Note that the concept of the constraint-based control has been recently applied to path planning or trajectory generation, e.g., in [38, 39, 40, 41, 42]. However, the above shortcomings of [34, 35, 36, 37] apply to the papers [38, 39, 40, 41], and all of [38, 39, 41, 42] focus on safety certificates and cannot be a solution to the area coverage problem.
@@
    We apply the constraint-based control to the high-level path generation, merging path generation and path following layers by exploiting a suitable trajectory parametrization for the nature of the robots considered. This can provide advantages because it combines the best of the two layers: coverage-performance informed decision making and real-time feedback.
@@ II Problem Formulation
The position coordinates of the origin and rotation angle of $\Sigma_{i}$ relative to the world frame $\Sigma_{\mathrm{w}}$ are denoted by $p_{i}\in\mathbb{R}^{2}$ and $\theta_{i}\in\mathbb{S}^{1}$, respectively. We denote the pose of USV $i$ relative to $\Sigma_{\mathrm{w}}$ as $z_{i}=[p_{i}^{\top}\ \theta_{i}]^{\top}$. We suppose that the motion of USV $i$ obeys the so-called Dubins vehicle model
$\dot{z}_{i}=\begin{bmatrix}\cos\theta_{i}&0\\ \sin\theta_{i}&0\\ 0&1\end{bmatrix}\nu_{i},\ \nu_{i}=\begin{bmatrix}\bar{v}\\ \omega_{i}\end{bmatrix},$ (1)
@@
where $\overline{\delta}$, $\underline{\delta}$ are positive scalars, and the operator
@@
In this article, we design an online path generator to efficiently sample data on all points $q_{j}\ (j=1,2,\dots,m)$. In principle, our solution can be applied to any path that meets the following conditions:
@@
Note that the present solution updates the path, and hence the actual trajectory that each USV follows can be more complex than the circular or elliptic path.
@@ II-A Circular Path Generation Problem
respectively. Solving the quadratic equation (5a) together with (5b) for $c_{i}$ yields two solutions, which correspond to the right-hand turning path, denoted by $\mathrm{r}$, and left-hand one, denoted by $\mathrm{l}$, as illustrated in Fig. 2. Thus, the turning direction becomes another decision variable for each USV $i$, which is denoted by ${\rm X}_{i}\in\mathcal{X}=\{\mathrm{r},\mathrm{l}\}$. Notice that once the turning direction ${\rm X}_{i}\in\mathcal{X}$ is fixed, we can eliminate the variable $c_{i}$ by solving (5) in advance. The center $c_{i}$ meeting (5) is given by
@@
following the manner of coverage control [11, 14, 15, 16], where $c=[c^{\top}_{1}\ c^{\top}_{2}\ \cdots\ c^{\top}_{n}]^{\top}$, $r=[r_{1}\ r_{2}\ \cdots\ r_{n}]^{\top}$, and $\mathrm{X}=(\mathrm{X}_{i})_{i=1}^{n}$. Now, since both the function $J_{\rm c}^{\mathrm{X}}$ and the constraints (5) are changing in time due to the variations of $\phi_{j}(t)$ and $z_{i}(t)$ in real time, it is difficult to maximize $J_{\rm c}^{\mathrm{X}}$ while meeting (5) instantaneously. Also, the gradient-based methodology in [14] does not provide any performance guarantee. We thus relax the objective and consider the constraint-based specification
@@ III-B Elliptic Coverage Path Generator
In order to limit the size of the ellipse, we constrain the eigenvalues of the matrix $S_{i}$ as
@@
Especially for large USV fleets, there might be cases where no observation point $q_{j}$ is assigned to a USV as a result of partitioning in (28) or (41). When this happens, all the coefficients for the variable $\rho_{i}$ in the inequality constraint on the performance guarantee in (33) or (46) get equal to zero, and hence the constraint gets independent of the selection of $\rho_{i}$. Noticing that the cost in the QP is given by $\|\rho_{i}\|^{2}$ plus the cost for the slack variable, the optimal solution must be $\rho_{i}=0$ and hence the previous path parameters must be maintained. Thus, even when no observation point is assigned to a USV, no serious problem happens except for the violation of the local performance constraint.
@@
The Voronoi-like partition may have instantaneous big changes, and accordingly, the paths may significantly change, especially for large USV fleets. Even in the presence of such changes, the paths must continue to meet (5) or (15) as well as the size constraints (27) or (40). Consequently, such changes do not impose extreme actions on the USVs. In other words, the effect of possible instability of the Voronoi-based allocation is inherently mitigated by these constraints.
@@ IV Simulation with Ideal Mathematical Model
Figure 7: Time evolution of the functions $b^{2}_{\mathrm{e},1}$ (cyan), $b^{3}_{\mathrm{e},1}$ (orange), $b^{4}_{\mathrm{e},1}$ (yellow), and $b^{5}_{\mathrm{e},1}$ (purple) associated with the elliptic shape and size constraints for USV 1.
The movie of the simulation is uploaded to https://youtu.be/DfWsMAWdaZc, whose snapshots are shown in Fig. 5. The white curves show the trajectories of the USVs. Since the paths are updated online, the trajectories are no longer the ellipses. The color map of the environment shows the value of the importance index $\phi_{j}$, where red region indicates high values and blue low values. We observe from the movie and the snapshots that both USVs update the elliptic paths so that they pass through the red area.
The time evolution of the functions $b^{1}_{\mathrm{e},1}$ (red) and $b^{1}_{\mathrm{e},2}$ (blue) are shown in Fig. 7. We observe from the figure that $b^{1}_{\mathrm{e},i}\geq 0$ is satisfied at almost all times. This implies that the present path generator almost certifies the performance guarantee. However, the constraint is violated only at around $160$\mathrm{s}$$. In order to further investigate the violation, we show the time evolution of the functions $b^{k}_{\mathrm{e},1}\ (k=2,3,4,5)$ for USV 1 in Fig. 7. We see from this figure that $b^{5}_{\mathrm{e},1}$ is close to the lower limit, and the constraint associated with $b^{5}_{\mathrm{e},1}$ gets active during the period. We conclude from these figures that the constraint violations in Fig. 7 are caused by the relaxation of the performance constraints to soft constraints in (46). Note that the constraint on $b^{5}_{\mathrm{e},1}$ is not only violated by the initial state but also slightly violated at around 160s. The latter violation was caused by the unavoidable discretization in the path computation. It is, however, well-known that the constraint-based control inherently drives the states to satisfy the constraints even if the state gets out of the safe set for some reason [50, 51]. This is why these temporal constraint violations do not cause any serious problem.
From these results, we conclude that the present path generator works as expected, at least for an ideal mathematical model (1). Finally, in order to show that the proposed path generator works even for a larger fleet of USVs, a simulation video of the circular path generator with 6 USVs is uploaded to https://youtu.be/DOW4wdwdVX8 for reference.
@@ V-B Modeling and Local Controller Design
Figure 12: Time series data of the model outputs (blue), the actual output data (gray), and the input data (black).
@@
where the gains were tuned so that the gain crossover frequency was equal to 4.0 rad/s. A bigger crossover frequency was prohibited due to the actuator saturation.  [^4] The actuators suffer from saturation in order to prevent the thrusters from rotating backward. Specifically, both $u_{i}^{\rm right}$ and $u_{i}^{\rm left}$ in (50) should remain greater than 1500. The time response for the step reference $\omega_{i}^{\rm ref}=0.5$\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$$ is shown in Fig. 13. We see that $\omega_{i}$ successfully tracks to the reference $\omega_{i}^{\rm ref}$ despite the disturbances from the water.
@@ V-C Collision Avoidance against Walls
Refer to caption (a) $t=5$\mathrm{s}$$ Refer to caption (b) $t=30$\mathrm{s}$$ Refer to caption (c) $t=60$\mathrm{s}$$ Refer to caption (d) $t=120$\mathrm{s}$$ Refer to caption (e) $t=180$\mathrm{s}$$ Refer to caption (f) $t=240$\mathrm{s}$$ Refer to caption
Figure 15: Snapshots of the experiment at each time. The left side shows moving two real USVs on the experiment pool. The red and green circular paths on the right side indicate paths corresponding to USVs 1 and 2. White curves show the trajectories of the two USVs after $t=0$\mathrm{s}$$.
@@ V-D Experimental Verification
We implemented the proposed control architecture for two USVs under the settings above. The movie of the experiment is uploaded to https://youtu.be/YuZbbcMRrLU, whose snapshots are shown in Fig. 15. We see from the movie and figures that the USVs switch the turning directions $\mathrm{X}_{i}$ and radius $r_{i}$ so that they are persistently driven to the red regions with high importance indices, while avoiding overlaps of the monitoring areas between the two USVs.
@@
Let us finally show the actual monitoring performance achieved by the USVs. To this end, we show in Fig. 18 the time evolution of the sum of the overall importance indices, $\sum_{j=1}^{m}\phi_{j}$, which is regarded as a reasonable metric for the monitoring performance. It is seen from this figure that it initially decays sharply and then is maintained below $2.1\times 10^{3}$. The fact that the performance metric stays at a certain value means that each observation point has been monitored by the USVs persistently. It is thus concluded that the present coverage path generator successfully achieves persistent environmental monitoring even in the presence of uncertainties in the physical world. Note, however, that the relation between the performance level $\gamma$ for the path generation and the actual environmental monitoring performance remains implicit. A direct evaluation of the latter is left as future work.
@@ V-E Comparative Evaluation with Baseline Algorithm
These paths are discretized by $0.2$\mathrm{m}$$, and the waypoints $\varpi_{i}^{1},\varpi_{i}^{2},\cdots,\varpi_{i}^{\ell}$ are assigned at the endpoints of the path segments, with $\varpi_{i}^{\ell+1}=\varpi_{i}^{1}$. The closest segment is then selected as the initial target path segment for a path following controller based on the Line-of-Sight guidance algorithm [30]. Given the target path segment $\varpi_{i}^{k}\varpi_{i}^{k+1}$, the reference angle for USV $i$ is determined as follows:
@@
The video of the baseline algorithm experiment is available at https://youtu.be/ACtlfoO1MlU. The path is set up so that USVs do not collide with the pool wall. This is especially the case when the USV follows a path segment connecting the end and beginning points of the path generated by the DCAC algorithm. It takes about $55$\mathrm{s}$$ for a USV to complete one lap. When the USV follows the straight line part, the observation area tends to cover more points with higher importance indices. When it follows the curves, it tends to cover points that have been recently observed. The comparison of the sum of importance indices between the baseline algorithm and the proposed method is shown in Fig. 20. In both cases, the USVs pass through unvisited points at the beginning, so that there is no significant difference in the first $10$\mathrm{s}$$. The steep declines continue in the first $50$\mathrm{s}$$ in both methods, but the coverage path generation achieves lower values than the lawnmower pattern. After the rapid decline, the values are stabilized, albeit with some oscillations. The results demonstrate that our proposed method achieves more efficient coverage by reducing the sum of importance indices more effectively than the baseline algorithm.
In addition, the present method is advantageous over the conventional path planning methods in the following aspects. First, the proposed path generator explicitly reflects the decay/increase rate of the importance indices which correspond to the required amount of the sensing data and the required timing of resampling. Second, the generator can compensate for the coverage holes caused, e.g., by disturbances due to the real-time feedback of the current importance indices. Third, the generator has flexibility in adding or deleting USVs in the middle of the operation, since the area partitions are updated in real time. Remark finally that the baseline algorithm covers the upper-left and lower-right regions, differently from the present path generator. This is because the baseline algorithm does not consider the wall-avoidance constraints. If visiting all points is of importance, randomly switching the priority of left-side and right-side avoidance over time allows USVs to achieve this even in the presence of the wall-avoidance constraints.